OpenStudy (gold_n_diamonds):

4 radians is the same as _____?

OpenStudy (steve816):

\[\large 4 ~radians = \frac{ 4 * 180 }{ \pi} ~degrees\]

OpenStudy (gold_n_diamonds):

Thanks

OpenStudy (steve816):

No problem! Welcome to OpenStudy!!!

OpenStudy (shamim):

anyway do u know pi radians=180 degree

OpenStudy (kittiwitti1):

As @shamim and @steve816 said, the conversion factor for radians to degrees is\[x\pi\text{ radians}=x\pi\times\frac{180°}{\pi}\]\[\rightarrow x\cancel{\pi}^{1}\times\frac{180°}{\cancel{\pi}^{1}}\rightarrow x\times180°=y°\] Don't forget to close this when you are done :-)

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)